Two projectiles, P1 and P2, are fired at different angles from a common starting line and ultimately land the same distance away on level ground. If P1 is launched at 25.0° from horizontal, what can be concluded about P2's angle of launch from horizontal, time in air, and speed at which it lands compared to P1?


A) launch angle= 45.0°, time in air= less than P1, landing speed= same as P1

B) launch angle= 45.0°, time in air= same as P1, landing speed= zero

C) launch angle= 65.0°, time in air= greater than P1, landing speed= same as P1

D) launch angle= 65.0°, time in air= less than P1, landing speed= greater than P1


1
Expert's answer
2021-04-19T07:22:06-0400

The horizontal distance covered by a projectile is given as follows:


"d = \dfrac{v^2\sin^2\theta}{2g}"

where "v" is the launch speed (which is equal to landing speed), "\theta" is the launch angle, "g" is the gravitational acceleration. Since the distance is the same for both P1 and P2, obtain:

"\dfrac{v_1^2\sin^2\theta_1}{2g} = \dfrac{v_2^2\sin^2\theta_2}{2g}\\
v_1^2\sin^2\theta_1 = v_2^2\sin^2\theta_2"

Expressing the landing speed of P2, obtain:


"v_2 = v_1\dfrac{\sin\theta_1}{\sin\theta_2}"

The launch angle of P1 is "\theta_1 = 25\degree". Suppose "\theta_2 = 45\degree\space or\space \theta_2 = 65\degree". In any case ratio "\dfrac{\sin\theta_1}{\sin\theta_2}>1" and the speed of P1 is greater than the speed of P2.


Answer. D.